# Beyond Isolation: Multi-Agent Synergy for Improving Knowledge Graph Construction

### v2

@@ 4.1. Construction of Multi-expert Agent Collaboration Network
In-context Demonstration. Some studies (Wan et al., 2023; Ma et al., 2023; Liu et al., 2022b) show improvements in contextual learning by selecting few-shot demonstrations based on similarity. Our contextual prompts $\mathcal{P}_{c}$ are introduced as N-way K-shot sampling of the demonstration samples $\mathcal{M}=N\times K$, providing direct evidence about the task and references to predictions. However, limited by the input tokens of the LLMs, a single prompt may not contain all supported instances, so we use a sentence embedding similarity-based approach to select the $\mathcal{M}$ examples with the closest Euclidean distance as contexts.
4.3. Collaborative Interaction Optimization
For devising intelligent agent optimization methods, ChatIE (Wei et al., 2023) introduces an innovative approach centered on task decomposition and the selection of pertinent extraction objects through two-stage artificial templates. Departing from the conventional strategy of devising fixed-stage thinking rounds tailored to a singular task, we reach to the periphery of the age-old adage, ”Two heads are better than one.” In the context of team collaboration optimization, the need for meticulous decomposition design diminishes, as proficient agents specializing in diverse domains engage in reflective interactions, honing their responses organically. As shown in Algorithm 1, after collecting replicas by other expert agents, we further provide collaboration prompts $\mathcal{P}_{v}$: The relation extraction answer you gave in the last round of collaboration was "##LAST_ROUND_RESULT##". The answer given by the NER expert agent in the knowledge graph construction team was "##NER_RESULT##". The EE expert agent gave The answer that comes out is "##EE_RESULT##". You should refer to the answers of other team members to add or delete some of your answers, and give credible reasons. If no modification is needed, please copy the answers from the previous round." Following ample interactions, the ultimate result materializes through the filtration of responses that deviate from the prescribed output format constraints (Section 3.1).
5. Experiments
  • RQ3: What are the benefits of replicas delivery simplification and customized expert knowledge background?
  • RQ4: How effective is the proposed CooperKGC in extracting different types of entities, relations and events?
5.1. Experiment Settings
5.1.1. Dataset.
As to the NER task, we conduct experiments on the following popular benchmark: Conllpp (Wang et al., 2019), OntoNotes5.0 (Pradhan et al., 2013) and MSRA (Levow, 2006). The Conllpp dataset is a modified version of the Conll2003 (Sang and Meulder, 2003) and contains 4 entity types. OntoNotes5.0 is an English NER dataset containing 18 types of named entities. MSRA dataset is a Chinese named entity recognition dataset for the news field and contains 3 entity types. For RE task, we conduct experiments on the following popular benchmark: NYT11-HRL (Takanobu et al., 2019), Re-TACRED (Stoica et al., 2021), and DuIE2.0 (Li et al., 2019). NYT11-HRL is a preprocessed version of NYT11 (Hoffmann et al., 2011) and contains 12 pre-defined relation types. Re-TACRED is a notably enhanced version of the TACRED dataset (Zhang et al., 2017) for relation extraction which contains over 91 thousand sentences spread across 40 relations. DuIE2.0 is the industry’s largest schema-based Chinese RE dataset and contains 48 predefined relation types. Some of the objects in the triples have multiple attributes, called complex-object values. There are two standard EE datasets: ACE05 (Walker et al., 2006) and DuEE1.0 (Li et al., 2020). The ACE05 dataset provides event annotations in document and sentence levels from a variety of domains by defining 33 event types and 35 role types. DuEE1.0 dataset is a Chinese event extraction dataset released by Baidu, which contains 65 event types.
Figure 3. Results of equipping KGC agents with different expert knowledge backgrounds.
5.1.2. Baselines.
In our experimental framework, we opt for AutoKG (Zhu et al., 2023) as the implementation of Vanilla LLMs for KGC realm, which defines an end-to-end extraction workflow through the manual templates. Expanding on this foundation, ChatIE (Wei et al., 2023) refines the extraction process using a two-round method. Taking RE as an example, this method entails the initial extraction of the relation, followed by the output of the associated entity span. This sequential approach mirrors a cognitive model’s thought process, explicitly delineating the steps of task decomposition. Further, CoT-ER (Ma et al., 2023) introduces an explicit evidence reasoning method, characterized by three rounds of processing. In the first and second rounds, the LLM is required to output concept-level entities corresponding to head and tail entities. Subsequently, in the third round, the extraction of relevant entity spans occurs, establishing a specific relationship between these two entities with explicit evidence.
5.1.3. Evaluation Protocols.
For the NER task, results is assessed using micro F1, with the predicted entity’s span should align precisely with the groundtruth, and the entity type should be an accurate match. Shifting to the RE task, the standard micro F1 metric is adopted for assessment. Correctness is contingent upon the accuracy of both the subject and object entity spans, coupled with the correct identification of the relation type. For the EE task, micro F1 is applied on the ACE05 dataset, necessitating the predicted event triggers and arguments to precisely align with the groundtruth. While on the DuEE1.0 dataset, the F1-measure relies on word-level matching, ensuring consistency in identification and classification of event triggers and arguments.
5.2. Performance Comparison with SOTA (RQ1)
@@
In our study, we present comprehensive results from both 0-shot and 1-shot experiments conducted across 8 datasets. Note that due to API cost considerations, each experiment is conducted with 100 samples from the test/valid set and the results are evaluated using the standard micro F1. To facilitate fair comparisons, we adopt the ”gpt-3.5-turbo” API as the default LLM backend for both baseline models and our proposed methods. The ”temperature” parameter is set to 0.0, indicating the stable output of GPT-3.5, while we report the average results of three runs for all experiments. Within our method, the maximum of rounds to 4, and the number of KGC team members is fixed to 3. For the English dataset, the default customized expert knowledge background for the NER task is based on Conllpp, the RE task is NYT11-HRL, and the EE task is ACE05. Similarly, for the Chinese dataset, the NER task is based on MSRA, the RE task is DUIE2.0, and the EE task is DUIE1.0. Taking the OntoNotes5.0 dataset evaluation as an example, our team members’ expert knowledge background is informed by the specifications of OntoNotes5.0, NYT11-HRL, and ACE05. This consistent methodology ensures a coherent and standardized approach across diverse datasets and tasks.
In Table 2, we report the F1-score results for 3 KGC tasks on the 8 datasets. Observations are as follows:
@@
(1) CooperKGC improves the overall performance in diverse tasks. As can be seen from the NER task results, CooperKGC outperforms the vanilla LLM extraction method (AutoKG) in the 0-shot setting, where on Conllp dataset improved by 10.7, OntoNotes5.0 improved by 13.4, and MSRA improved by 3.4. Similarly, CooperKGC shows a general improvement in the 1-shot setting, which indicates that our method achieves significant improvements in both zero-shot and few-shot learning scenarios. In addition, CooperKGC is significantly higher than the baseline on all 8 datasets and 3 types of tasks, illustrating the advantages of the collaborative architecture. As compared to a single execution, the teamwork prompt strategy shows the benefits of associating different information extraction tasks. We believe that part of the improvement can be attributed to the multipath architecture based on extraction information sharing, which allows for an objective view of the opinions expressed by different experts. In contrast, for a simple extraction approach like AutoKG with a single round of LLM calls, on the one hand, the overly heavy information input for task comprehension and rule constraints poses a challenge for a single model. On the other hand there is a lack of sufficient inference steps for a self-debugging process. n our approach, multiple rounds of interactions alleviate this anxiety of requiring ”hit-and-miss” reasoning, making it easier to explicitly identify erroneous intermediate feedbacks during the interactions.
@@
(2) Teamwork is an effective implicit reasoning chain. Taking the NYT11-HRL dataset as an example, although ChatIE improved by 25.0 over the baseline in the 0-shot setting while achieving an improvement of 18.2 over the baseline in the 1-shot setting, we believe that the gain stems from decomposing the extraction process into two phases. Among the first stage is determining the types of relations involved in a given sentence, which often involves multiple relations in a single sentence. The second stage then designs triple extraction templates for each relation, and similarly a relation can involve multiple triples. This is a kind of explicit expression of ”Let’s Think Step by Step(Kojima et al., 2022), which clearly indicates the sub-tasks to be accomplished in each stage of LLM, and thus achieves improved results. Further CoT-ER uses side information to induce LLM to generate explicit evidence of relationship reasoning, resulting in an improvement of 32.8 over the baseline in the 0-shot setting, as well as an improvement of 20.9 over the baseline in the 1-shot setting. Intuitively, triple extraction can be viewed as a mapping of head-to-tail entity spans to relation-type reasoning. Therefore, we believe that the gain comes from more intensive bootstrapping of LLM-specific task-level and concept-level knowledge. CooperKGC outperforms both, with a 33.2 improvement in the 0-shot and a 22.7 improvement in the 1-shot setting. We believe that building collaborative teams contributes to ”Brain Storming(Osborn, 1953), where each round of the brainstorming process is performed by the members of team. By collecting evidence from other members in each round of interactions, agent’s responses is fine-tuned from the previous round. Although there is no reasoning path planned for LLM, this proactive optimisation shows more encouraging prospects than passive methods.
5.3. Analysis of Team Members and Interaction Rounds (RQ2)
Table 3. Performance under different member assignments.
@@ 5.3. Analysis of Team Members and Interaction Rounds (RQ2)
Figure 4. Performance with Increased Rounds.
To further investigate the impact brought by the combination of intelligences with different expert knowledge backgrounds on team collaboration, we introduce an experiment to analyse the diverse combination of team members. Specifically, We experiment on 0-shot setting and the number of team members is fixed to 3. By replacing the expert knowledge backgrounds representing NER agent, RE agent, and EE agent, we analyse which kind of expert knowledge backgrounds (mainly the schema constraints in the task description $\mathcal{P}_{t}$) could produce better benefits for the team construction goals. Figure 3 shows the results of equipping KGC agents with different expert knowledge backgrounds, and we observe that combination b (OntoNotes5.0+RE-TACRED+ACE05) allows EE expert agents to achieve the best extraction performance, and the richer variety of relation types guided by RE-TACRED allows EE agents to discover more potential arguments compared with combination a. In addition, combination b achieves a more comprehensive improvement compared to combination c. We analyse the schema of OntoNotes5.0 versus Conllp and find that three of the entity categories are the same (”PER”, ”LOC”, ”ORG”), while the remaining 15 more specialised entity categories refine the ”MISC” category in Conllp, which results in benefits in extraction performance for the RE-TACRED and ACE05 datasets. We therefore conclude that more specialised expert agents, i.e., equipped with fine-grained schema constraints, can bring more insightful information to guide teamwork.
Figure 5. Ablation study of CooperKGC framework.
Table 4. A qualitative example of CooperKGC on expert agent collaboration. Each expert absorbs a more detailed interactive view from different perspectives, leading to stronger knowledge selection, knowledge integration, and knowledge teamwork. Red text represents elements that conflict with groundtruth, and blue text represents elements that are updated compared to previous round’s result.
@@
0
(LOC, Palestinian section of the border crossing)
(Palestinian section of the border crossing, location-located_in, Israeli troops) (Israeli troops, person-nationality, Palestinians)
@@
[{Trigger Type: Conflict-Attack, Trigger Word: taken over, Arguments: (Attacker, Israeli troops),(Target: Palestinian section of the border crossing)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, the Palestinian section of the border crossing)}]
1
(PER, Israeli troops), (LOC, border), (ORG, police)
(Israeli troops, location-located_in, the Palestinian section of the border crossing), (Israeli troops, person-nationality, Israeli), (Six Palestinian police officers, person-nationality, Palestinians)
[{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, the Palestinian section of the border crossing)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, the Palestinian section of the border crossing)}]
2
(PER, Israeli troops), (LOC, border), (ORG, police), (PER, Six Palestinian police officers)
(Israeli troops, person-place_lived, the Palestinian section of the border crossing), (Israeli troops, person-nationality, Israeli), (Six Palestinian police officers, person-nationality, Palestinians)
[{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, Israeli)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, border), (Artifact, Israeli troops)}]
3
(PER, Israeli troops), (LOC, border), (ORG, police), (PER, Six Palestinian police officers)
(Israeli troops, person-place_lived, the Palestinian section of the border crossing), (Six Palestinian police officers, person-nationality, Palestinians)
[{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, Israeli)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, border), (Artifact, Six Palestinian police officers)}]
4
(PER, Israeli troops), (LOC, border), (ORG, police), (PER, Six Palestinian police officers)
(Israeli troops, person-place_lived, the Palestinian section of the border crossing), (Six Palestinian police officers, person-nationality, Palestinians)
@@
[{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, Israeli)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, the Palestinian section of the border crossing), (Artifact, Six Palestinian police officers)}]
@@
Another question is whether it is possible to equip with more agents to make more gains for our team. Table 3 shows the results of both kinds of experiments, the upper one is to add additional agents to the original team, and the results show that Team (3-Agent+BOTH) makes the extraction results of ACE05 improved by adding a NER Agent and a RE Agent. However, another risk is also demonstrated, in both Team (3-Agent+OntoNotes) and Team (3-Agent+RE-TACRED) it is observed that when more authoritative expert agents are introduced, it leads to a decrease in the extraction results of the agent for the same task, and this kind of unconscious opinion conformity is consistent with the concept of ”Presentation of Self(Goffman et al., 2002) in sociology. In addition, inspired by ”Self-consistency(Wang et al., 2023c), in the bottom of Table 3 we explore the difference between the performance of the self-consistent voting method and CooperKGC on a single task. Although the consistency method to some degree mitigates the randomness of the single agent producing the hallucinatory fact, it is nevertheless weaker than our results on all 3 representative datasets. We argue that a single perspective is unable to access the interactive information provided by other experts, and thus suffers from ”Information Cocoons(Sunstein, 2006).
@@
5.4. Ablation Study of CooperKGC (RQ3)
To further validate the rationality of the components, we undertakes a comprehensive examination through two ablation experiments: ”w/o Replicas Delivery Simplification Function” and ”w/o Customized Expert Knowledge Background.” As shown in Figure 5, the removal of simplification function led to a noticeable degradation in performance across all three datasets. We believe that besides bringing the advantage of reducing the length of token input, this filtered information method enhances the robustness of intra-team decision-making, avoids the noise distraction of individual thought processes, and makes team collaboration more concise and efficient. In the second experiment, we delved into the consequences of removing the tailored expert knowledge background. Although we retained plain schema guidance for each task, the lack of a self-cognitive prompt causes agents to be unskilled in locating the information from the text, consistent with research (Li et al., 2023a; Xu et al., 2023; Deshpande et al., 2023; Fu et al., 2023) that assigning roles can significantly affect their generative behaviour. We argue that agents with different specialisations can collaborate efficiently with each other in performing complex tasks. Furthermore, This role-playing mechanism allows multi-agents to interact actively with each other for their own task goals, alleviating the monotonous confidence in their own responses.
5.5. Case Study of Collaboration Process (RQ4)
To illustrate the effectiveness of our proposed CooperKGC for collaborative interactions in KGC teams, Table 4 provides a qualitative example demonstrating the intermediate process. Note the CoT reflection process such as ”After considering the extraction results of other agents…” is skipped, and the input sentence is an example of EE task ”Six Palestinian police officers were allowed to return to the Palestinian section of the border crossing, which had been taken over by Israeli troops shortly after the start of the uprising. ” we compare the results of the EE agent with the groundtruth, while the results of the NER agent and the RE agent are only for reference since there is no groundtruth. The observations are as follows: (1) Knowledge Selection. In Round 2, the EE agent borrows the LOC entity ”border” newly discovered by the NER agent in the previous round and adds an argument (Destination, border) to the original answer; (2) Knowledge Correction. In the 1st round of interactions, the EE agent corrects the wrong trigger word ”taken over”, which indicates that the team members have the ability to provide self-feedback; (3) Knowledge Aggregation. Although the EE agent puts a wrong argument (Destination, border) in round 3, it rectifies the hallucination facts generated in the interim by eliciting LLM semantic comprehension during the interaction.
6. Conclusion and Future Work
In this work, we have made an initial attempt to aggregate agents with different expert knowledge to form KGC team. Our findings reveal the impressive collaborative capabilities of LLM agents, demonstrating the possibilities of agent networks in mutually enhancing task performance. The human-like behaviours that emerge during collaboration resonate with sociological theories and effectively improve LLM performance in factuality, knowledge integration and intellectual reasoning. In the future, sociologically derived architectures could provide insightful inspirations. The rich variants of CooperKGC can be applied to multi-subject cognitive teams to solve more flexible collaborative tasks. In addition, as LLM evolves, understanding the adaptive combination patterns of team members will be critical to instructing more collaborative agent networks.

### v3

@@ 4.2 Performance Comparison with SOTA (RQ1)
Our study conducts comprehensive experiments on 0-shot and 1-shot settings across 8 datasets, each with 100 samples from test/valid sets, evaluating results using micro F1. We use the "gpt-3.5-turbo" API for both baseline models and proposed methods, with a temperature parameter set to 0.0 and average results reported over three runs. Our method sets a maximum of 4 rounds and 3 KGC team members. For the English dataset, the default customized expert knowledge background for the NER task is based on Conllpp, the RE task is NYT11-HRL, and the EE task is ACE05. Similarly, for the Chinese dataset, the NER task is based on MSRA, the RE task is DUIE2.0, and the EE task is DUIE1.0. Table 2 presents F1-score results for 3 KGC tasks across datasets, revealing:
@@
(1) CooperKGC enhances overall performance across diverse tasks. Compared to the vanilla method, CooperKGC shows significant improvements in both 0-shot and 1-shot settings. In contrast, for a simple extraction approach like AutoKG with a single round of LLM calls, on the one hand, the overly heavy information input for task comprehension and rule constraints poses a challenge for a single model. On the other hand there is a lack of sufficient inference steps for a self-debugging process. Our approach, multiple rounds of interactions alleviate this anxiety of requiring "hit-and-miss" reasoning, making it easier to explicitly identify erroneous intermediate feedbacks during the interactions.
@@
(2) Teamwork is an effective implicit reasoning chain. Taking the NYT11-HRL dataset as an example, although ChatIE improved by 25.0 over the baseline in the 0-shot setting while achieving an improvement of 18.2 over the baseline in the 1-shot setting, we believe that the gain stems from decomposing the extraction process into two phases. Among the first stage is determining the types of relations involved in a given sentence, which often involves multiple relations in a single sentence. The second stage then designs triple extraction templates for each relation, which clearly indicates the sub-tasks to be accomplished in each stage. CoT-ER uses head-to-tail mapping to induce LLM to generate explicit evidence of reasoning, resulting in an improvement of 32.8 over the baseline in the 0-shot setting. CooperKGC outperforms both, with a 33.2 improvement in the 0-shot and a 22.7 improvement in the 1-shot setting. We believe that building collaborative teams contributes to "Brain Storming" [18], where each round of the brainstorming process is performed by the members of team. By collecting evidence from other members in each round of interactions, agent’s responses is fine-tuned from the previous round. Although there is no reasoning path, this proactive optimisation shows more encouraging prospects than passive methods.
4.3 Analysis of Team Members and Interaction Rounds (RQ2)
Figure 2: Equipping KGC agents with different expert knowledge backgrounds.
To further investigate the impact brought by the combination of intelligences with different expert knowledge backgrounds on team collaboration, we introduce an experiment to analyse the diverse combination of team members. Specifically, We experiment on 0-shot setting and the number of team members is fixed to 3. By replacing the expert knowledge backgrounds representing NER agent, RE agent, and EE agent, we analyse which kind of expert knowledge backgrounds (mainly the schema constraints in the task description $\mathcal{P}_{t}$) could produce better benefits for the team construction goals. Figure 2 shows the results of equipping KGC agents with different expert knowledge backgrounds, and we observe that combination b (OntoNotes5.0+RE-TACRED+ACE05) allows EE expert agents to achieve the best extraction performance, and the richer variety of relation types guided by RE-TACRED allows EE agents to discover more potential arguments compared with combination a. In addition, combination b achieves a more comprehensive improvement compared to combination c. We analyse the schema of OntoNotes5.0 versus Conllp and find that three of the entity categories are the same ("PER", "LOC", "ORG"), while the remaining 15 more specialised entity categories refine the "MISC" category in Conllp, which results in benefits in extraction performance for the RE-TACRED and ACE05 datasets. We therefore conclude that more specialised expert agents, i.e., equipped with fine-grained schema constraints, can bring more insightful information to guide teamwork.
Table 3: Micro-F1 Performance under different member assignments.
@@ 4.3 Analysis of Team Members and Interaction Rounds (RQ2)
Figure 3: Micro-F1 Performance with Increased Rounds.
Table 4: A qualitative example of CooperKGC on expert agent collaboration.
@@
0 (LOC, Palestinian section of the border crossing) (Palestinian section of the border crossing, location-located_in, Israeli troops) (Israeli troops, person-nationality, Palestinians) [{Trigger Type: Conflict-Attack, Trigger Word: taken over, Arguments: (Attacker, Israeli troops),(Target: Palestinian section of the border crossing)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, the Palestinian section of the border crossing)}]
@@
1 (PER, Israeli troops), (LOC, border), (ORG, police) (Israeli troops, location-located_in, the Palestinian section of the border crossing), (Israeli troops, person-nationality, Israeli), (Six Palestinian police officers, person-nationality, Palestinians) [{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, the Palestinian section of the border crossing)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, the Palestinian section of the border crossing)}]
2 (PER, Israeli troops), (LOC, border), (ORG, police), (PER, Six Palestinian police officers) (Israeli troops, person-place_lived, the Palestinian section of the border crossing), (Israeli troops, person-nationality, Israeli), (Six Palestinian police officers, person-nationality, Palestinians) [{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, Israeli)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, border), (Artifact, Israeli troops)}]
3 (PER, Israeli troops), (LOC, border), (ORG, police), (PER, Six Palestinian police officers) (Israeli troops, person-place_lived, the Palestinian section of the border crossing), (Six Palestinian police officers, person-nationality, Palestinians) [{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, Israeli)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, border), (Artifact, Six Palestinian police officers)}]
4 (PER, Israeli troops), (LOC, border), (ORG, police), (PER, Six Palestinian police officers) (Israeli troops, person-place_lived, the Palestinian section of the border crossing), (Six Palestinian police officers, person-nationality, Palestinians) [{Trigger Type: Conflict-Attack, Trigger Word: uprising, Arguments: (Attacker, Israeli troops), (Place, Israeli)}, {Trigger Type: Movement:Transport, Trigger Word: return, Arguments: (Destination, the Palestinian section of the border crossing), (Artifact, Six Palestinian police officers)}]
@@
Another question is whether it is possible to equip with more agents to make more gains for our team. Table 3 shows the results of both kinds of experiments, the upper one is to add additional agents to the original team, and the results show that Team (3-Agent+BOTH) makes the extraction results of ACE05 improved by adding a NER Agent and a RE Agent. However, another risk is also demonstrated, in both Team (3-Agent+OntoNotes) and Team (3-Agent+RE-TACRED) it is observed that when more authoritative expert agents are introduced, it leads to a decrease in the extraction results of the agent for the same task, and this kind of unconscious opinion conformity is consistent with the concept of "Presentation of Self" [6] in sociology. In addition, inspired by "Self-consistency" [29], in the bottom of Table 3 we explore the difference between the performance of the self-consistent voting method and CooperKGC on a single task. Although the consistency method to some degree mitigates the randomness of the single agent producing the hallucinatory fact, it is nevertheless weaker than our results on all 3 representative datasets. We argue that a single perspective is unable to access the interactive information provided by other experts, and thus suffers from "Information Cocoons"[22].
@@
4.4 Case Study of Collaboration Process (RQ3)
To illustrate the effectiveness of our proposed CooperKGC for collaborative interactions in KGC teams, Table 4 provides a qualitative example demonstrating the intermediate process. Note the CoT reflection process such as "After considering the extraction results of other agents…" is skipped, and the input sentence is an example of EE task "Six Palestinian police officers were allowed to return to the Palestinian section of the border crossing, which had been taken over by Israeli troops shortly after the start of the uprising. " we compare the results of the EE agent with the groundtruth, while the results of the NER agent and the RE agent are only for reference since there is no groundtruth. The observations are as follows: (1) Knowledge Selection. In Round 2, the EE agent borrows the LOC entity "border" newly discovered by the NER agent in the previous round and adds an argument (Destination, border) to the original answer; (2) Knowledge Correction. In the 1st round of interactions, the EE agent corrects the wrong trigger word "taken over", which indicates that the team members have the ability to provide self-feedback; (3) Knowledge Aggregation. Although the EE agent puts a wrong argument (Destination, border) in round 3, it rectifies the hallucination facts generated in the interim by eliciting LLM semantic comprehension during the interaction.
5 Conclusion and Future Work
In this study, we initiated the formation of a KGC team by aggregating agents with diverse expertise. Our results highlight the collaborative potential of LLM agents, showcasing how agent networks can enhance task performance collectively. The emergence of human-like behaviors in collaboration aligns with sociological theories, leading to improvements in factuality, knowledge integration, and intellectual reasoning. Future research could draw insights from sociologically derived architectures, expanding the application of CooperKGC variants to solve diverse collaborative tasks.